# DeNOTS: Stable Deep Neural ODEs for Time Series

### v4

@@ B.1 Time scaling
$\left|\frac{d}{dt}\|\Delta\|_{2}\right|=\left|\frac{d}{dt}\sqrt{\Delta\cdot\Delta}\right|=\frac{\Delta\cdot\Delta^{\prime}}{\|\Delta\|_{2}}\leq\frac{\|\Delta\|_{2}\cdot\|\Delta^{\prime}\|_{2}}{\|\Delta\|_{2}}=\|\Delta^{\prime}\|_{2},$
where we denote $\Delta^{\prime}\triangleq\frac{d}{dt}\Delta$. Consequently, taking the norm of (11), we get:
$\|\mathbf{g}_{\theta}(\hat{\mathbf{x}}_{1}(t),\mathbf{h}_{1}(t))-\mathbf{g}_{\theta}(\hat{\mathbf{x}}_{2}(t),\mathbf{h}_{2}(t))\|_{2}=\|\Delta^{\prime}\|_{2}\geq\left|\frac{d}{dt}\|\Delta\|_{2}\right|\geq\frac{d}{dt}\|\Delta\|_{2}.$
Now we use the Lipschitzness of $\mathbf{g}_{\theta}$:
$\frac{d}{dt}\|\Delta\|_{2}\leq\|\hat{\mathbf{x}}_{1}(t)-\hat{\mathbf{x}}_{2}(t)\|_{2}M_{x}+\|\Delta\|_{2}M_{h}.$
Suppose that $\|\Delta(t)\|_{2}=C(t)e^{M_{h}t}$. Substituting, we get:
$\frac{d}{dt}\|\Delta\|_{2}=C^{\prime}e^{M_{h}t}+M_{h}Ce^{M_{h}t}\leq\|\hat{\mathbf{x}}_{1}(t)-\hat{\mathbf{x}}_{2}(t)\|_{2}M_{x}+CM_{h}e^{M_{h}t},$
where $C^{\prime}\triangleq\frac{dC}{dt}$. We isolate $dC$ to integrate the resulting expression:
$dC\leq\|\hat{\mathbf{x}}_{1}(t)-\hat{\mathbf{x}}_{2}(t)\|_{2}M_{x}e^{-M_{h}t}dt\implies\int_{t=0}^{T}dC\leq\int_{t=0}^{T}\|\hat{\mathbf{x}}_{1}(t)-\hat{\mathbf{x}}_{2}(t)\|_{2}M_{x}e^{-M_{h}t}dt.$
Finally, we bound the integral on the right:
$\int_{t=0}^{T}dC=C(T)-C(0)$ $\leq\int_{t=0}^{T}\|\hat{\mathbf{x}}_{1}(t)-\hat{\mathbf{x}}_{2}(t)\|_{2}M_{x}e^{-M_{h}t}dt$
$\leq M_{x}\left(\int_{t=0}^{T}\|\hat{\mathbf{x}}_{1}(t)-\hat{\mathbf{x}}_{2}(t)\|_{2}dt\right)\left(\int_{t=0}^{T}e^{-M_{h}t}dt\right)$
$\leq\rho(\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2})M_{x}\frac{1}{M_{h}}(1-e^{-M_{h}T}).$
Since $\|\Delta(0)\|_{2}=0$, we have $C(0)=0$. Substituting into $\|\Delta(t)\|_{2}=C(t)e^{M_{h}t}$, we get the desired expression. ∎
@@ B.3 Stability
As mentioned in the main text, we approach stability analysis through control systems theory Sontag & Wang (1995). For clarity, we write out the definitions and theorems used in greater detail. First, we introduce the comparison function classes that are used in the stability definition:
@@
The connection between Definitions B.2 and B.3 is given by the following lemma Sontag & Wang (1995):
@@ B.4 Error Bounds
B.4.1 Setting
Let $\mathbf{x}(t)$ be a stationary zero-mean $u$-dimensional Gaussian process (GP) with the covariance function $\mathbf{K}(\cdot,\cdot)=\{K_{i}(\cdot,\cdot)\}$ and independent components:
@@ B.4.1 Setting
B.4.2 Robustness theorem
We will now estimate, how much of the variance from (8) accumulates in the final hidden state:
Theorem (4.5).
Robustness (Gaussian Process standpoint). Along with (3) under Assumptions 4.14.2 consider analogous ODE dynamics for a realization of the actual Gaussian Process $\mathbf{x}(t)$ (instead of its MAP approximation $\hat{\mathbf{x}}(t)$):
$\begin{cases}\frac{d\mathbf{h}^{*}}{dt}=a\mathbf{f}_{\theta}(\mathbf{x}(t),\mathbf{h}^{*}(t))-b\mathbf{h}^{*}(t),\\ \mathbf{h}^{*}(0)=\mathbf{h}(0)=\mathbf{h}_{0}.\end{cases}$
Then, depending on whether we are working with point-wise or interval errors, we can write different bounds on the variance, accumulating in $bh$.
For the pointwise bound:
$\mathbb{E}\lVert\mathbf{h}(t_{n})-\mathbf{h}^{*}(t_{n})\rVert_{2}^{2}\leq\left(\frac{\bar{\sigma}_{\mathrm{PW}}aL_{x}}{b-aL_{h}}\right)^{2}.$ (9)
For the interval bound:
$\frac{1}{n}\int_{0}^{T}\mathbb{E}\left\lVert\mathbf{h}(t)-\mathbf{h}^{*}(t)\right\rVert_{2}^{2}dt\leq\left(\frac{\bar{\sigma}_{\mathrm{INT}}aL_{x}}{b-aL_{h}}\right)^{2}.$ (10)
Proof.
$=\left(a(\mathbf{f}_{\theta}(\hat{\mathbf{x}}(t),\mathbf{h}(t))-\mathbf{f}_{\theta}(\mathbf{x}(t),\mathbf{h}^{*}(t)))-b(\mathbf{h}(t)-\mathbf{h}^{*}(t)),\mathbf{h}(t)-\mathbf{h}^{*}(t)\right)=$
$=a(\mathbf{f}_{\theta}(\hat{\mathbf{x}}(t),\mathbf{h}(t))-\mathbf{f}_{\theta}(\mathbf{x}(t),\mathbf{h}^{*}(t)),\delta\mathbf{h}(t))-b\|\delta\mathbf{h}(t)\|_{2}^{2}=$
$\leq a(L_{h}\left\lVert\delta\mathbf{h}\right\rVert_{2}^{2}+L_{x}\left\lVert\delta\mathbf{x}(t)\right\rVert_{2}\left\lVert\delta\mathbf{h}\right\rVert_{2})-b\left\lVert\delta\mathbf{h}\right\rVert_{2}^{2}=$
$=(aL_{h}-b)\left\lVert\delta\mathbf{h}\right\rVert_{2}^{2}+aL_{x}\left\lVert\delta\mathbf{x}(t)\right\rVert_{2}\left\lVert\delta\mathbf{h}\right\rVert_{2},$ (15)
where we introduce the following notation for conciseness:
$\delta\mathbf{h}(t)$ $\triangleq\mathbf{h}(t)-\mathbf{h}^{*}(t),$
$\delta\mathbf{x}(t)$ $\triangleq\hat{\mathbf{x}}(t)-\mathbf{x}(t).$
If we have a pointwise bound on $\mathbb{E}\left\lVert\delta\mathbf{x}(t)\right\rVert_{2}^{2}$, the expression from (15) can be upper bounded by a quadratic function, where the multiplier $aL_{h}-b$ next to the squared term is negative, since $aL_{h}<b$. It turns negative for larger $\left\lVert\delta\mathbf{h}(t)\right\rVert_{2}$:
$\mathrm{if}\;\;\left\lVert\delta\mathbf{h}\right\rVert_{2}>\frac{aL_{x}\bar{\sigma}_{\textrm{PW}}}{b-aL_{h}},\;\;\textrm{then}\;\;(aL_{h}-b)\left\lVert\delta\mathbf{h}(t)\right\rVert_{2}^{2}+L_{x}\sigma_{\textrm{PW}}\left\lVert\delta\mathbf{h}(t)\right\rVert_{2}<0$
which proves half of this theorem.
However, if we only have an interval error bound, we are forced to integrate (15) on the interval $t_{1}=0,t_{n}=T]$:
$\frac{1}{2}\int_{0}^{T}\frac{d}{dt}\left\lVert\mathbf{h}(t)-\mathbf{h}^{*}(t)\right\rVert_{2}^{2}dt\leq\int_{0}^{T}\left\lVert\delta\mathbf{h}(t)\right\rVert_{2}^{2}(aL_{h}-b)dt+\int_{0}^{T}L_{x}\left\lVert\delta\mathbf{x}(t)\right\rVert_{2}\left\lVert\delta\mathbf{h}(t)\right\rVert_{2}dt.$ (16)
$\int_{0}^{T}\left\lVert\delta\mathbf{x}(t)\right\rVert_{2}\left\lVert\delta\mathbf{h}(t)\right\rVert_{2}dt$ $\leq\sqrt{\int_{0}^{T}\left\lVert\delta\mathbf{x}(t)\right\rVert_{2}^{2}dt\int_{0}^{T}\left\lVert\delta\mathbf{h}(t)\right\rVert_{2}^{2}dt}\triangleq XH\sqrt{k},$
where we introduced the following notation:
$X^{2}\triangleq\int_{0}^{T}\left\lVert\delta\mathbf{x}(t)\right\rVert_{2}^{2}dt,\;\;H^{2}\triangleq\int_{0}^{T}\left\lVert\delta\mathbf{h}(t)\right\rVert_{2}^{2}dt.$
In light of that, let’s rewrite (16):
$\frac{1}{2}\left(\left\lVert\delta\mathbf{h}(T)\right\rVert_{2}^{2}-\left\lVert\delta\mathbf{h}(0)\right\rVert_{2}^{2}\right)=\frac{1}{2}\left\lVert\delta\mathbf{h}(T)\right\rVert_{2}^{2}\leq(aL_{h}-b)H^{2}+aL_{x}XH\sqrt{k}.$
The right-hand side is a quadratic function with respect to $H$; again, the multiplier next to the squared term is negative. Simultaneously, the left-hand side is positive, being a norm. Together, this means that $H$ needs to be in between the roots of this quadratic function to satisfy the positivity constraints:
$0\leq H\leq\frac{aL_{x}X\sqrt{k}}{b-aL_{h}}.$
We conclude the proof by taking the expectation of the square and factoring in the interval bound, which transforms into $\mathbb{E}X^{2}\leq\bar{\sigma}_{\textrm{INT}}^{2}$ due to linearity. ∎
B.4.3 Assumptions
@@
Prior work Zaytsev & Burnaev (2017) demonstrated that the results for infinite sequences empirically hold for finite ones. Indeed, since a typical kernel decays with an exponential speed, distant observations negligibly impact the resulting error.
B.4.4 Error bounds
@@
The optimal estimator for a uniform infinite grid is well-known Kolmogorov (1941):
@@
where optimal $\mathbf{K}$ is a symmetric kernel that depends on a spectral density. This form allows us to deduce analytical error bounds. For this purpose, we generalise the results from Zaytsev & Burnaev (2017) to multiple dimensions:
@@
After factoring out the spectral density integral, the terms (18), 19 and (20) are the expansion of a binomial:
@@
With Lemma B.10, all other results can be taken directly from Zaytsev & Burnaev (2017), taking into account that $\mathbf{x}(t)$ is $u$-dimensional. We recount select findings from this paper: the general error form, an analytic error expression for a specific kernel, and the minimax error bound.
@@
At the limit $\theta\rightarrow 0$, the mean of the resulting Gaussian Processes tends to a natural cubic spline Golubev & Krymova (2013). We estimate the corresponding interval error $\sigma_{\textrm{INT}}^{2}(\theta)$ at $\theta\rightarrow 0$, which can be interpreted as the interval error for natural cubic spline interpolation:
@@
where the spectral density $F$ is given by 23. Specifically, for splines, we are interested in the following limit:
@@
We devised a tight error bound for the interval error: indeed, the expression from Lemma 4.7 turns to equality for uniform sequences. It is also important to highlight, that pointwise bounds exist for the non-GP, deterministic setting (and can be plugged into (9), supposing that $\mathbf{x}$ is degenerate) Hall & Meyer (1976); De Boor & De Boor (1978), the optimal one being:
@@
B.4.5 Proof of special case of Assumption B.5.
@@
B.4.7 Bound Tightness
Example 1.
In this case, $b-aL_{h}=\varepsilon>0$. The solution is:
$h(t)=\frac{-Ae^{-\varepsilon t}+A}{\varepsilon}.$
$h^{*}(t)=\frac{-Be^{-\varepsilon t}+B}{\varepsilon}.$
B.5 Long-term Memory

### v5

@@ B.1 Time scaling
$\|\mathring{F}(\mathbf{x}_{1},\mathbf{h}_{0})-\mathring{F}(\mathbf{x}_{2},\mathbf{h}_{0})\|\triangleq\|\Delta\|\leq\rho(\mathbf{x}_{1},\mathbf{x}_{2})M_{x}\sqrt{\frac{1}{2M_{h}}\left(e^{2M_{h}t}-1\right)}$
Here, we will also denote the difference between these trajectories by $\Delta$:
$\mathring{F}(\mathbf{x}_{1},\mathbf{h}_{0})-\mathring{F}(\mathbf{x}_{2},\mathbf{h}_{0})=\Delta.$
We square the Lipschitz inequality, and take the derivative w.r.t. $t$:
$2(\Delta\mathbf{g},\Delta)\leq\frac{M_{x}^{2}}{2M_{h}}\left(\frac{d\rho^{2}}{dt}\left(e^{2M_{h}t}-1\right)+\rho^{2}\frac{d}{dt}\left(e^{2M_{h}t}-1\right)\right).$
Here, we use the notation $\Delta\mathbf{g}=\mathbf{g}(\mathbf{x}_{1}(t),\mathbf{h}_{1}(t))-\mathbf{g}(\mathbf{x}_{2}(t),\mathbf{h}_{2}(t)).$ The derivatives of $\rho^{2}$ and $\left(e^{2M_{h}t}-1\right)$ are easily calculated analytically:
$\frac{d\rho^{2}}{dt}=\|\mathbf{x}_{1}(t)-\mathbf{x}_{2}(t)\|^{2}.$
$\frac{d}{dt}\left(e^{2M_{h}t}-1\right)=2M_{h}e^{2M_{h}t}.$
We substitute these expressions:
$2(\Delta\mathbf{g},\Delta)\leq\frac{M_{x}^{2}}{2M_{h}}\left(\|\mathbf{x}_{1}(t)-\mathbf{x}_{2}(t)\|^{2}.\left(e^{2M_{h}t}-1\right)+\rho^{2}2M_{h}e^{2M_{h}t}.\right).$
Now we divide the above by $t$, and take the limit at $t\rightarrow 0$.
$\underset{t\rightarrow 0}{\lim}2(\Delta\mathbf{g},\frac{1}{t}\Delta)\leq\underset{t\rightarrow 0}{\lim}\frac{M_{x}^{2}}{2M_{h}}\left(\|\mathbf{x}_{1}(t)-\mathbf{x}_{2}(t)\|^{2}\frac{1}{t}\left(e^{2M_{h}t}-1\right)+\frac{1}{t}\rho^{2}2M_{h}e^{2M_{h}t}\right)$
We calculate the limits for terms with $\frac{1}{t}$:
$\underset{t\rightarrow 0}{\lim}\frac{1}{t}\Delta=\Delta\mathbf{g}(0);$
$\underset{t\rightarrow 0}{\lim}\frac{1}{t}\left(e^{2M_{h}t}-1\right)=2M_{h};$
$\underset{t\rightarrow 0}{\lim}\frac{1}{t}\rho^{2}=\|\mathbf{x}_{1}(0)-\mathbf{x}_{2}(0)\|^{2}.$
The limit for the other terms is achieved by simply setting $t=0$:
$2(\Delta\mathbf{g}(0),\Delta\mathbf{g}(0))\leq\frac{M_{x}^{2}}{2M_{h}}\left(\|\mathbf{x}_{1}(0)-\mathbf{x}_{2}(0)\|^{2}2M_{h}+\|\mathbf{x}_{1}(0)-\mathbf{x}_{2}(0)\|^{2}2M_{h}\right).$
Simplifying, we get:
$\|\mathbf{g}(\mathbf{x}_{1}(0),\mathbf{h}_{0})-\mathbf{g}(\mathbf{x}_{2}(0),\mathbf{h}_{0})\|^{2}\leq M_{x}\|\mathbf{x}_{1}(0)-\mathbf{x}_{2}(0)\|^{2}.$
Since the choice of $\mathbf{x}_{1}(0),\mathbf{x}_{2}(0),\mathbf{h}_{0}$ is arbitrary, the above implies that $\mathbf{g}$ is $M_{x}$-Lipschitz.
Now for the $h$ component
. We follow a very similar approach, but this time the input trajectories are the same $\mathbf{x}_{1}=\mathbf{x}_{2}=\mathbf{x}$, while the starting points are not $\mathbf{h}_{1}\neq\mathbf{h}_{2}$. Therefore, since $F$ is Lipschitz, we can write the following inequality:
$\|\mathring{F}(\mathbf{x},\mathbf{h}_{1};t)-\mathring{F}(\mathbf{x},\mathbf{h}_{2};t)\|\triangleq\|\Delta(t)\|\leq\|\Delta_{0}\|\left(e^{M_{h}t}-1\right).$
Here, $\Delta_{0}\triangleq\mathbf{h}_{1}-\mathbf{h}_{2}$. Again, we take the square and differentiate w.r.t. $t$:
$2(\Delta,\Delta\mathbf{g})\leq\|\Delta_{0}\|^{2}2\left(e^{M_{h}t}-1\right)M_{h}e^{M_{h}t}.$
Divide by $t$ and take the limit at $t\rightarrow 0$:
$2\|\Delta\mathbf{g}\|^{2}\leq\|\Delta_{0}\|^{2}2M_{h}.$
Since the choice of $\mathbf{x},\mathbf{h}_{1},\mathbf{h}_{2}$ is arbitrary, we have proven that $\mathbf{g}$ is $M_{h}$-Lipschitz.
This concludes the proof of Theorem 3.1. ∎
As a corollary, given an $M_{x},M_{h}$-Lipschitz NN $\mathbf{g}$, we can represent all the $L_{x},L_{h}$-Lipschitz (differentially-definable, start-corrected) $F$, and only them.
@@ B.3 Stability
As mentioned in the main text, we approach stability analysis through control systems theory sontag1995characterizations. For clarity, we write out the definitions and theorems used in greater detail. First, we introduce the comparison function classes that are used in the stability definition:
@@
The connection between Definitions B.2 and B.3 is given by the following lemma sontag1995characterizations:
@@ B.4 Error Bounds
To devise the theoretical results of this section, we need to introduce the ”perfect“ Neural CDE, calculated without any errors:
$\mathbf{h}^{*}(0)=\mathbf{h}_{0};\quad\frac{d\mathbf{h}^{*}(t)}{dt}=a\mathbf{f}_{\theta}(\mathbf{x}(t),\mathbf{h}^{*}(t))-b\mathbf{h}^{*}(t).$ (6)
Then, both the integration and the interpolation errors may be modelled via the following ”imperfect“ ODE, often referred to as the ”Modified Vector Field“ reich1999backward:
$\mathbf{h}(0)=\mathbf{h}_{0};\quad\frac{d\mathbf{h}(t)}{dt}=a\mathbf{f}_{\theta}(\mathbf{x}(t),\mathbf{h}(t))-b\mathbf{h}(t)+\xi(t),$ (7)
where $\xi(t)$ is a random vector, which represents the momentary deviation of our numerical ODE from the true one. For further theoretical considerations, we introduce two types of bounds on $\xi(t)$:
$\text{Pointwise: }\mathbb{E}\left\lVert\xi(t)\right\rVert_{2}^{2}\leq\xi_{\mathrm{PW}}^{2},\forall t;$ $\text{Interval: }\frac{1}{t_{b}-t_{a}}\int_{t_{a}}^{t_{b}}\mathbb{E}\left\lVert\xi(t)\right\rVert_{2}^{2}dt\leq\xi_{\mathrm{INT}}^{2},\forall t_{a}<t_{b}.$ (8)
For each of the introduced bounds, we can devise an estimate for the error in the final hidden state:
Theorem (4.5).
Error Bounds (Robustness). Consider the differential equations (6),(7), under Assumptions 4.14.2 Then, depending on whether we are working with the pointwise or the interval errors (8), we can write different bounds on the variance, accumulating in $\mathbf{h}$.
Pointwise bound: $\mathbb{E}\lVert\mathbf{h}(T)-\mathbf{h}^{*}(T)\rVert_{2}^{2}\leq\left(\frac{a}{b-aL_{h}}\right)^{2}\xi_{\mathrm{PW}}^{2}.$ (9)
Interval bound: $\frac{1}{T}\int_{t_{a}}^{t_{b}}\mathbb{E}\left\lVert\mathbf{h}(t)-\mathbf{h}^{*}(t)\right\rVert_{2}^{2}dt\leq\left(\frac{a}{b-aL_{h}}\right)^{2}\xi^{2}_{\mathrm{INT}}.$ (10)
Proof.
$=\left(a\delta_{f}-b\delta_{h}+\xi,\delta_{h}\right)=$
$=a(\delta_{f}+\xi,\delta_{h})-b\|\delta_{h}\|_{2}^{2}=$
$\leq a(L_{h}\left\lVert\delta_{h}\right\rVert_{2}^{2}+\left\lVert\xi\right\rVert\left\lVert\delta_{h}\right\rVert_{2})-b\left\lVert\delta_{h}\right\rVert_{2}^{2}=$
$=(aL_{h}-b)\left\lVert\delta_{h}\right\rVert_{2}^{2}+a\left\lVert\xi\right\rVert_{2}\left\lVert\delta_{h}\right\rVert_{2},$ (16)
where we introduce the following notation for conciseness:
$\delta_{f}$ $\triangleq\mathbf{f}_{\theta}(\mathbf{x},\mathbf{h})-\mathbf{f}_{\theta}(\mathbf{x},\mathbf{h}^{*})$
$\delta_{h}$ $\triangleq\mathbf{h}(t)-\mathbf{h}^{*}(t),$
$\delta_{x}$ $\triangleq\hat{\mathbf{x}}(t)-\mathbf{x}(t).$
If we have a pointwise bound on $\mathbb{E}\left\lVert\xi\right\rVert_{2}^{2}$, the expression from (16) can be upper bounded by a quadratic function, where the multiplier $aL_{h}-b$ next to the squared term is negative, since $aL_{h}<b$. It turns negative for larger $\left\lVert\delta_{h}\right\rVert_{2}$:
$\mathrm{if}\;\;\left\lVert\delta_{h}\right\rVert_{2}>\frac{a\xi_{\textrm{PW}}}{b-aL_{h}},\;\;\textrm{then}\;\;(aL_{h}-b)\left\lVert\delta_{h}\right\rVert_{2}^{2}+\xi_{\textrm{PW}}\left\lVert\delta_{h}\right\rVert_{2}<0$
which proves half of this theorem.
However, if we only have an interval error bound, we are forced to integrate (16) on the interval $[0,T]$:
$\frac{1}{2}\int_{0}^{T}\frac{d}{dt}\left\lVert\mathbf{h}(t)-\mathbf{h}^{*}(t)\right\rVert_{2}^{2}dt\leq\int_{0}^{T}\left\lVert\delta_{h}(t)\right\rVert_{2}^{2}(aL_{h}-b)dt+\int_{0}^{T}\left\lVert\xi(t)\right\rVert_{2}\left\lVert\delta_{h}(t)\right\rVert_{2}dt.$ (17)
$\int_{0}^{T}\left\lVert\xi(t)\right\rVert_{2}\left\lVert\delta_{h}(t)\right\rVert_{2}dt$ $\leq\sqrt{\int_{0}^{T}\left\lVert\xi(t)\right\rVert_{2}^{2}dt\int_{0}^{T}\left\lVert\delta_{h}(t)\right\rVert_{2}^{2}dt}\triangleq RH,$
where we introduced the following notation:
$R^{2}\triangleq\int_{0}^{T}\left\lVert\xi(t)\right\rVert_{2}^{2}dt,\;\;H^{2}\triangleq\int_{0}^{T}\left\lVert\delta_{h}(t)\right\rVert_{2}^{2}dt.$
In light of that, let’s rewrite (17):
$\frac{1}{2}\left(\left\lVert\delta_{h}(T)\right\rVert_{2}^{2}-\left\lVert\delta_{h}(0)\right\rVert_{2}^{2}\right)=\frac{1}{2}\left\lVert\delta_{h}(T)\right\rVert_{2}^{2}\leq(aL_{h}-b)H^{2}+aRH\sqrt{k},$
where we used $\delta_{h}(0)=0$. The right-hand side is a quadratic function with respect to $H$; again, the multiplier next to the squared term is negative. Simultaneously, the left-hand side is positive, being a norm. Together, this means that $H$ needs to be in between the roots of this quadratic function to satisfy the positivity constraints:
$0\leq H\leq\frac{aR\sqrt{k}}{b-aL_{h}}.$
We square the above, and then substitute $R,H$ in:
$\int_{0}^{T}\left\lVert\delta_{h}(t)\right\rVert_{2}^{2}\leq\left(\frac{a}{b-aL_{h}}\right)^{2}\int_{0}^{T}\left\lVert\xi(t)\right\rVert_{2}^{2}dt.$
Dividing by $T$ and taking the expectation (which we may move under the integral due to linearity), we get:
$\frac{1}{T}\int_{0}^{T}\mathbb{E}\left\lVert\delta_{h}(t)\right\rVert_{2}^{2}\leq\left(\frac{a}{b-aL_{h}}\right)^{2}\frac{1}{T}\int_{0}^{T}\mathbb{E}\left\lVert\xi(t)\right\rVert_{2}^{2}dt\leq\left(\frac{a}{b-aL_{h}}\right)^{2}\xi_{\textrm{INT}}.$
B.4.1 Bound Tightness
Example 1.
In this case, $b-aL_{h}=\varepsilon>0$. The solution is:
$h(t)=\frac{-Ae^{-\varepsilon t}+A}{\varepsilon}.$
$h^{*}(t)=\frac{-Be^{-\varepsilon t}+B}{\varepsilon}.$
B.5 GP Interpolation Error
Now, let’s devise the Gaussian Process interval-error estimates. Let $\mathbf{x}(t)$ be a stationary zero-mean $u$-dimensional Gaussian process (GP) with the covariance function $\mathbf{K}(\cdot,\cdot)=\{K_{i}(\cdot,\cdot)\}$ and independent components:
@@ B.5 GP Interpolation Error
B.5.1 GP Assumptions
@@
Prior work zaytsev2017minimax demonstrated that the results for infinite sequences empirically hold for finite ones. Indeed, since a typical kernel decays with an exponential speed, distant observations negligibly impact the resulting error.
B.5.2 Interpolation Error bounds
@@
The optimal estimator for a uniform infinite grid is well-known kolmogorov1941interpolation:
@@
where optimal $\mathbf{K}$ is a symmetric kernel that depends on a spectral density. This form allows us to deduce analytical error bounds. For this purpose, we generalise the results from zaytsev2017minimax to multiple dimensions:
@@
After factoring out the spectral density integral, the terms (23), 24 and (25) are the expansion of a binomial:
@@
With Lemma B.10, all other results can be taken directly from zaytsev2017minimax, taking into account that $\mathbf{x}(t)$ is $u$-dimensional. We recount select findings from this paper: the general error form, an analytic error expression for a specific kernel, and the minimax error bound.
@@
At the limit $\theta\rightarrow 0$, the mean of the resulting Gaussian Processes tends to a natural cubic spline golubev2013interpolation. We estimate the corresponding interval error $\sigma_{\textrm{INT}}^{2}(\theta)$ at $\theta\rightarrow 0$, which can be interpreted as the interval error for natural cubic spline interpolation:
@@
where the spectral density $F$ is given by 28. Specifically, for splines, we are interested in the following limit:
@@
We devised a tight error bound for the interval error: indeed, the expression from Lemma 4.6 turns to equality for uniform sequences. It is also important to highlight, that pointwise bounds exist for the non-GP, deterministic setting (and can be plugged into (9), supposing that $\mathbf{x}$ is degenerate) hall1976optimal; de1978practical, the optimal one being:
@@
B.5.3 Proof of special case of Assumption B.5.
@@
B.6 Long-term Memory